ituedrgeiru:

Simplify the expression. one half cubed − 6 ÷ square root of 64 negative five eighths negative 47 over 64 five eighths 47 over 64

ituedrgeiru:

answer choices negative five eighths negative 47 over 64 five eighths 47 over 64

Joe348:

\[1/2^3=0.125\] \[0.125-6=-5.875\] \[\sqrt{64} =8\] \[-5.875 \div 8\] \[=-0.73...\] As a fraction \[-47/64\]

Joe348:

So B!